OpenStudy (anonymous):

In how many ways can a student work 7 out of 10 questions on an exam?

OpenStudy (anonymous):

i am going to guess that there was more to the problem for example, the test is true/false

OpenStudy (anonymous):

unless it is just this: a student must work out 7 out of 10 questions on an exam. how many ways can she select the 7 questions?

OpenStudy (anonymous):

Yes...that's all the questions asked.

OpenStudy (anonymous):

you copied it exactly?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

maybe written by someone for whom english is a third language

OpenStudy (anonymous):

8) In how many ways can a student work 7 out of 10 questions on an exam? 8) A) 10,000,000 ways B) 720 ways C) 21 ways D) 120 ways

OpenStudy (anonymous):

i am going to guess that you are being asked to compute \[\dbinom{10}{7}\] which you might also see as \(_{10}C_7\)

OpenStudy (anonymous):

ah ok yes, that is what you want \[_{10}C_7=\frac{10\times 9\times 8}{3\times 2}=10\times 3\times 4=120\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw